# Crepant Transformation correspondence For Toric Stack Bundles

### v1

@@ 2.2.1. Equivariant Cohomology of Toric Stacks
[$[\left\{(z_{1},\dots,z_{m})\in U_{\omega}:z_{i}=0\right\}/K],$
@@
$\mathfrak{I}=\left\langle\chi-\sum_{i=1}^{m}\langle\chi,b_{i}\rangle u_{i}:\chi\in\mathbf{N}^{\vee}\otimes\mathbb{C}\right\rangle,\quad\mathfrak{J}=\left\langle\prod_{i\notin I}u_{i}:I\notin\mathcal{A}_{\omega}\right\rangle$
@@
The $T$-equivariant cohomology ring of $X_{\omega}$ is given by the extension of scalars
$H^{\bullet}_{T}(X_{\omega};\mathbb{C})\cong H^{\bullet}_{\mathcal{Q}}(X_{\omega};\mathbb{C})\otimes_{R_{\mathcal{Q}}}R_{T}$
@@ 2.3. Wall and Chamber Structure
The space of stability conditions $\omega\in\mathbb{L}^{\vee}\otimes\mathbb{R}$ has a wall and chamber structure, which describes how the GIT quotient $X_{\omega}$ changes as $\omega$ varies.
@@
The quotient $X_{\omega^{\prime}}$ is isomorphic to $X_{\omega}$ for any $\omega^{\prime}\in C_{\omega}$. However it changes when $\omega^{\prime}$ crosses a codimension one boundary (called a wall) of $C_{\omega}$.
@@
Let $C_{+}$ and $C_{-}$ be two chambers in $\mathbb{L}^{\vee}\otimes\mathbb{R}$ separated by a hyperplane $W$. Pick stability conditions $\omega_{+}\in C_{+},\omega_{-}\in C_{-}$. Consider $\mathcal{A}_{\pm}=\mathcal{A}_{\omega_{\pm}}$ as defined above, and the corresponding toric stacks $X_{\pm}=X_{\omega_{\pm}}$. Let $\varphi:X_{+}\dashrightarrow X_{-}$ be the birational transformation induced by the toric wall-crossing. We make the additional assumption
@@
so that transformation is crepant. Let $e\in W^{\perp}\cap\mathbb{L}$ be a primitive lattice vector such that $e$ is positive on $C_{+}$ and negative in $C_{-}$.
@@
$\mathbb{K}_{\pm}=\left\{f\in\mathbb{L}\otimes\mathbb{Q}:I_{f}\in\mathcal{A}_{\pm}\right\}$
where $I_{f}=\{i\in[m]:D_{i}\cdot f\in\mathbb{Z}\}$. Let $\tilde{\mathbb{L}}_{\pm}$ be the free $\mathbb{Z}$-submodule generated by $\mathbb{K}_{\pm}$. Note that $D_{j}$ is always in $\tilde{\mathbb{L}}_{\pm}^{\vee}$ for any $j\in S_{\pm}$ and they are linearly independent, so we may extend $\{D_{j}:j\in S_{\pm}\}$ to a basis of $\tilde{\mathbb{L}}_{\pm}^{\vee}$. After reordering, the basis can be written as $\left\{p^{\pm}_{1},\dots,p^{\pm}_{r-1},p^{\pm}_{r}\right\}$ such that $p^{+}_{i}=p^{-}_{i}\in W$ for $11\leq i\leq r-$ 1 and $p^{\pm}_{r}$ is the unique vector in each basis that does not lie on $W$. Let $\{y^{\pm}_{1},\dots,y^{\pm}_{r}\}$ be the corresponding coordinates, and define
$(y^{\pm})^{d}=\prod_{i=1}^{r}(y^{\pm}_{i})^{p^{\pm}_{i}\cdot d}$
for any $d\in\mathbb{L}\otimes\mathbb{R}$. Under change of coordinates, we have $(y^{+})^{e}=(y^{-})^{e}$.
@@
One can interpret $\mathcal{M}$ as the moduli space of certain Landau-Ginzburg models (mirror to the toric stacks). For this paper, our use of $\mathcal{M}$ is mainly $y^{\pm}$, which are variables in $I$-functions. See [8] for a detailed discussion about $\mathcal{M}$.
@@ 2.4. Toric Stack Bundles
Let $P\rightarrow B$ be a principal (×$(\mathbb{C}^{\times})^{m}$-bundle over a smooth projective variety $B$. Let $K,D_{i}$ and $U_{\omega}$ be as in Section 2.2. Associated to these data, we define a toric stack bundle to be the quotient
@@
where $K$ acts on $U_{\omega}$ via characters $D_{i}$ and acts on $P$ trivially. More details about construction of toric stack bundles can be found in [13].
@@
$\text{Box}(\mathbf{\Sigma}_{\omega})=\left\{v\in\mathbf{N}:\overline{v}=\sum_{i\notin I}c_{i}\overline{b_{i}}\in\mathbf{N}\otimes\mathbb{R}\ \text{for some}\ I\in\mathcal{A},0\leq c_{i}<1\right\}$
@@
$[f]\leftrightarrow v_{f}=\sum_{i=1}^{m}\lceil-(D_{i}\cdot f)\rceil b_{i}.$
@@
where the age $\iota_{f}=\sum_{i\notin I_{f}}\langle D_{i}\cdot f\rangle$.
Now assume $P$ arises from a direct sum of line bundles $\oplus_{j=1}^{m}L_{j}$ by removing the zero section, and let $\Lambda_{j}=c_{1}(L_{j}^{*})$. We may take $K_{\mathbb{R}}=(S^{1})^{r}$ and regard the quotient (2.2) as a fiberwise symplectic reduction for the action of $K_{\mathbb{R}}$. This defines $r$ tautological bundles over $\mathcal{P}$. Define $-P_{1},\dots,-P_{r}$ to be the $T$-equivariant first Chern classes of the tautological bundles, and $p_{1},\dots,p_{r}$ the restrictions of $P_{1},...,P_{r}$ to the fiber $X_{\omega}$. Recall that $\lambda_{1},\dots,\lambda_{m}$ are the first Chern classes of the $m$ duals of tautological bundles on $BT$ and $R_{T}:=H^{\bullet}_{T}(\text{pt};\mathbb{C})=\mathbb{C}[\lambda_{1},\dots,\lambda_{m}]$. The classes $P_{1},\dots,P_{r}$ generate the cohomology ring $H^{\bullet}_{T}(\mathcal{P})$ over $H^{\bullet}(B)\otimes R_{T}$ just as $p_{1},\dots,p_{r}$ generate $H^{\bullet}_{T}(X_{\omega})$ over $R_{T}$. This gives a basis of $H^{\bullet}_{T}(\mathcal{P})$ over $H^{\bullet}(B)\otimes R_{T}$ consisting of monomials in $\{P_{i}\}$.
@@ 2.4.1. Wall-Crossing for Toric Stack Bundles
By construction of toric stack bundles (2.2), as the stability parameter $\omega$ crosses a wall, $\mathcal{P}$ changes as the fiber $X_{\omega}$ changes.
As in Section 2.3, consider chambers $C_{\pm}\subset\mathbb{L}\otimes\mathbb{R}$ separated by a wall $W$. Let $\omega_{\pm}\in C_{\pm}$ be stability conditions and
@@
${\tilde{\mathcal{P}}}$${\mathcal{P}_{+}}$${\mathcal{P}_{-}}$${\mathcal{P}_{0}}$$\scriptstyle{f_{+}}$$\scriptstyle{f_{-}}$$\scriptstyle{\varphi}$$\scriptstyle{g_{+}}$$\scriptstyle{g_{-}}$
where each map is induced from map on toric fibers. Especially, the canonical bundles $K_{\mathcal{P}_{\pm}}$ of $\mathcal{P}_{\pm}$ are restrictions of the same line bundle $L_{0}$ on $\mathcal{P}_{0}$ given by a line bundle on $B$ and the character $-\sum_{i=1}^{m}D_{i}$. The condition $\sum_{i=1}^{m}D_{i}\in W$ ensures that $L_{0}$ comes from a $\mathbb{Q}$-Cartier divisor $\mathcal{D}_{0}$ on $\mathcal{P}_{0}$. Then we have
$f^{*}_{+}(K_{\mathcal{P}_{+}})=f^{*}_{+}\circ g^{*}_{+}(\mathcal{D}_{0})=f^{*}_{-}\circ g^{*}_{-}(\mathcal{D}_{0})=f^{*}_{-}(K_{\mathcal{P}_{-}})$
@@ 3. Crepant Transformation Correspondence
In this Section we prove Theorem 0.1. First we compute the analytic continuation of the $I$-function using the Mellin-Barnes method which gives the linear transformation $\mathbb{U}$. Then we compute the Fourier-Mukai transformation on $K$-groups and compare it with $\mathbb{U}$.
@@ 3.1. $I$-Function and $H$-Function
The $I$-function for a toric stack bundle $\mathcal{P}$ is defined [14] to be
(3.1) $I^{\mathcal{P}}(y,z)=ze^{\sigma/z}\sum_{\mathfrak{D}\in\overline{NE}(B)}J^{B}_{\mathfrak{D}}(z,\tau)Q^{\mathfrak{D}}\sum_{d\in\mathbb{K}}y^{d}\Delta\mathbf{1}_{\left[-d\right]}$
@@
The main result of [14] states that $I^{\mathcal{P}}$ lies on the Lagrangian cone for $\mathcal{P}$. Hence $I^{\mathcal{P}}$ can be used to determine genus 00 Gromov-Witten theory of $\mathcal{P}$. See [4] for the case of toric bundles.
@@
$={z^{-(\sum_{i=1}^{m}(D_{j}\cdot d+\langle D_{j}\cdot-d\rangle)+(\sum_{i=1}^{m}\Lambda_{j})\mathfrak{D})}}$
$=\frac{1}{z^{\sum_{i=1}^{m}D_{j}\cdot d-c_{1}(P)\mathfrak{D}}}\cdot\frac{1}{z^{\iota_{[-d]}}},$
@@
and $\rho=\theta(D_{1}+\cdots+D_{m})$. The notation deg is the degree operator for Chen-Ruan orbifold cohomology and $\text{deg}_{0}$ is the degree operator for ordinary cohomology, i.e. the degree without age shifting. The degrees for variables is given by $\text{deg}(z)=2$ and
@@ 3.2. Analytic Continuation of $H$-function via Localization
For a minimal anticone $\delta\in\mathcal{A}$, there is a corresponding $T$-fixed point in each fiber $X_{\omega}$ over $b\in B$, and together they form a gerbe $B_{\delta}$ over $B$. Let $i_{\delta}:B_{\delta}\rightarrow\mathcal{P}$ be the inclusion map. Similarly, for inertia stack $I\mathcal{P}$, a $T$-fixed gerbe corresponds to a pair $(\delta,f)$ where $\delta\in\mathcal{A}$ and $f\in\mathbb{K}/\mathbb{L}$ satisfying $D_{i}\cdot f\in\mathbb{Z}$ for all $i\in\delta$. Let $i_{(\delta,f)}:B_{(\delta,f)}\rightarrow\mathcal{P}_{f}$ be the inclusion to $I\mathcal{P}$. We have the following restriction of the $H$-function to a fixed gerbe:
@@
For the rest of this Section, any notation with index $+$ (respectively --) means it is for $\mathcal{P}_{+}$ (respectively $\mathcal{P}_{-}$). Especially, let $I_{\pm}$ denote their $I$-functions and $H^{\mathfrak{D}}_{\pm}$ denote their $H$-functions respectively. $(\delta_{+},f_{+})$ indexes a $T$-fixed gerbe in $I\mathcal{P}_{+}$ and $(\delta_{-},f_{-})$ indexes a $T$-fixed gerbe in $I\mathcal{P}_{-}$. By $\delta_{+}|\delta_{-}$ we mean $\delta_{+}=\{j_{1},\dots,j_{r-1},j_{+}\},\delta_{-}=\{j_{1},\dots,j_{r-1},j_{-}\}$ with $D_{j_{1}},\dots,D_{j_{r-1}}\in W$, $D_{j_{+}}\cdot e>0,D_{j_{-}}\cdot e<0$. By $(\delta_{+},f_{+})|(\delta_{-},f_{-})$ we mean $\delta_{+}|\delta_{-}$ and there exists $\alpha\in\mathbb{Q}$ such that $f_{-}=f_{+}+\alpha e$.
@@
$i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}=i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{-}$
@@
$i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}=\sum_{(\delta_{-},f_{-})|(\delta_{+},f_{+})}C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}i^{*}_{(\delta_{-},f_{-})}H^{\mathfrak{D}}_{-}$
@@
$C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}=e^{\frac{\pi iw}{D_{j_{-}}\cdot e}(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}(f_{+}-f_{-}))}\frac{\sin\pi(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}(f_{+}-f_{-}))}{(-D_{j_{-}}\cdot e)\sin\frac{\pi}{-D_{j_{-}}\cdot e}(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}(f_{+}-f_{-}))}$
$\prod_{j:D_{j}\cdot e<0,j\neq j_{-}}\frac{\sin\pi(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot f_{+})}{\sin\pi(\frac{U_{j}(\delta_{-})}{2\pi i}+D_{j}\cdot f_{-})},$
@@
Let $\delta_{-}\in\mathcal{A}_{-}$ and $\delta_{+}\in\mathcal{A}_{+}$ be minimal anticones such that $\delta_{-}|\delta_{+}$, and $j_{-}\in\delta_{-}$ the unique element which is not in $\delta_{+}$. Then for any $j\in[m]$ we have
$U_{j}(\delta_{+})=U_{j}(\delta_{-})+\frac{D_{j}\cdot e}{D_{j_{-}}\cdot e}U_{j_{-}}(\delta_{+})$
@@
$U_{j}-\lambda_{j}=c_{1}(U_{1}-\lambda_{1})+\cdots+c_{r-1}(U_{r-1}-\lambda_{r-1})+c_{j_{-}}(U_{j_{-}}-\lambda_{j_{-}})$
@@
$U_{j}(\delta_{-})-\lambda_{j}=\sum_{i=1}^{r-1}c_{i}(-\lambda_{i})+c_{j_{-}}(-\lambda_{j_{-}})$
$U_{j}(\delta_{+})-\lambda_{j}=\sum_{i=1}^{r-1}c_{i}(-\lambda_{i})+c_{j_{-}}(U_{j_{-}}(\delta_{+})-\lambda_{j_{-}})$
@@
The above proof shows this lemma can be generalized to $i^{*}_{\delta_{+}}\theta_{+}(p)=i^{*}_{\delta_{-}}\theta_{-}(p)+\frac{p\cdot e}{D_{j_{-}}\cdot e}U_{j_{-}}(\delta_{+})$ for any $p\in\mathbb{L}\otimes\mathbb{C}$. Especially, we have $\sigma_{+}(\delta)=\sigma_{-}(\delta)$ for $\delta\in\mathcal{A}_{-}\cap\mathcal{A}_{+}$ and $\sigma_{+}(\delta_{+})=\sigma_{-}(\delta_{-})+\frac{(\text{log}(y^{+}))^{e}}{D_{j_{-}}\cdot e}U_{j_{-}}(\delta_{+})$ for $\delta_{+}|\delta_{-}$.
@@
Define $\delta^{\vee}_{+}=\{d\in\mathbb{L}\otimes\mathbb{R}:D_{j}\cdot d+\Lambda_{j}\mathfrak{D}\geq 0,\forall j\in\delta_{+}\}$, note that the summation in $i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}$ is actually over $d\in\delta^{\vee}_{+}$. Also for $j\in\delta_{+}$, $D_{j}\cdot e=0$ except for one $j_{+}$ where $D_{j_{+}}\cdot e>0$, so for any $d\in\delta^{\vee}_{+}$ we may write $d=d_{+}+ke$ where $d_{+}\in\delta^{\vee}_{+}$ but $d_{+}-e\notin\delta^{\vee}_{+}$, $k\in\mathbb{Z}_{\geq 0}$. Then

### v2

@@ 2.2.1. Equivariant Cohomology of Toric Stacks
$\left\{(z_{1},\dots,z_{m})\in U_{\omega}:z_{i}=0\right\}/K],$
@@
$\mathfrak{I}=\left\langle\chi-\sum_{i=1}^{m}\langle\chi,b_{i}\rangle u_{i}:\chi\in\mathbf{N}^{\vee}\otimes\mathbb{C}\right\rangle,\quad\mathfrak{J}=\left\langle\prod_{i\notin I}u_{i}:I\notin\mathcal{A}_{\omega}\right\rangle.$
@@
The $T$-equivariant cohomology ring of $X_{\omega}$ is given by the extension of scalars,
$H^{\bullet}_{T}(X_{\omega};\mathbb{C})\cong H^{\bullet}_{\mathcal{Q}}(X_{\omega};\mathbb{C})\otimes_{R_{\mathcal{Q}}}R_{T},$
@@ 2.3. Wall and Chamber Structure
The space of stability conditions $\mathbb{L}^{\vee}\otimes\mathbb{R}$ has a wall and chamber structure, which describes how the GIT quotient $X_{\omega}$ changes as $\omega\in\mathbb{L}^{\vee}\otimes\mathbb{R}$ varies.
@@
The quotient $X_{\omega^{\prime}}$ is isomorphic to $X_{\omega}$ for any $\omega^{\prime}\in C_{\omega}$. However $X_{\omega^{\prime}}$ changes when $\omega^{\prime}$ crosses a codimension one boundary (called a wall) of $C_{\omega}$.
@@
Let $C_{+}$ and $C_{-}$ be two chambers in $\mathbb{L}^{\vee}\otimes\mathbb{R}$ separated by a hyperplane $W$. Pick stability conditions $\omega_{+}\in C_{+},\omega_{-}\in C_{-}$. Consider $\mathcal{A}_{\pm}=\mathcal{A}_{\omega_{\pm}}$ as defined above, and the corresponding toric stacks $X_{\pm}=X_{\omega_{\pm}}$. Let $\varphi:X_{+}\dashrightarrow X_{-}$ be the birational transformation induced by the toric wall-crossing. We make the additional assumption:
@@
so that transformation $\varphi$ is crepant. Let $e\in W^{\perp}\cap\mathbb{L}$ be a primitive lattice vector such that $e$ is positive on $C_{+}$ and negative in $C_{-}$.
@@
$\mathbb{K}_{\pm}=\left\{f\in\mathbb{L}\otimes\mathbb{Q}:I_{f}\in\mathcal{A}_{\pm}\right\},$
where $I_{f}=\{i\in[m]:D_{i}\cdot f\in\mathbb{Z}\}$. Let $\tilde{\mathbb{L}}_{\pm}$ be the free $\mathbb{Z}$-submodule generated by $\mathbb{K}_{\pm}$. Note that $D_{j}$ is always in $\tilde{\mathbb{L}}_{\pm}^{\vee}$ for any $j\in S_{\pm}$ and they are linearly independent, so we may extend $\{D_{j}:j\in S_{\pm}\}$ to a basis of $\tilde{\mathbb{L}}_{\pm}^{\vee}$. After reordering, the basis can be written as $\left\{p^{\pm}_{1},\dots,p^{\pm}_{r-1},p^{\pm}_{r}\right\}$ such that $p^{+}_{i}=p^{-}_{i}\in W$ for $1\leq i\leq r-1$ and $p^{\pm}_{r}$ is the unique vector in each basis that does not lie on $W$. Let $\{y^{\pm}_{1},\dots,y^{\pm}_{r}\}$ be the corresponding coordinates, and, for any $d\in\mathbb{L}\otimes\mathbb{R}$, define
$(y^{\pm})^{d}=\prod_{i=1}^{r}(y^{\pm}_{i})^{p^{\pm}_{i}\cdot d}.$
Under change of coordinates, we have $(y^{+})^{e}=(y^{-})^{e}$.
@@
One can interpret $\mathcal{M}$ as the moduli space of certain Landau-Ginzburg models (mirror to the toric stacks). For this paper, our use of $\mathcal{M}$ is mainly to house $y^{\pm}$, which are variables in $I$-functions. See [8] for a detailed discussion about $\mathcal{M}$.
@@ 2.4. Toric Stack Bundles
Let $P\rightarrow B$ be a principal $(\mathbb{C}^{\times})^{m}$-bundle over a smooth projective variety $B$. Let $K,D_{i}$ and $U_{\omega}$ be as in Section 2.2. Associated to these data, we define a toric stack bundle to be the quotient
@@
where $K$ acts on $U_{\omega}$ via characters $D_{i}$ and acts on $P$ trivially. $\mathcal{P}$ is a fiber bundle with fiber $X_{w}$ over $B$ and it is constructed from the principal $(\mathbb{C}^{\times})^{m}$-bundle $P\to B$ by changing fibers. More details about construction of toric stack bundles can be found in [13].
@@
$\text{Box}(\mathbf{\Sigma}_{\omega})=\left\{v\in\mathbf{N}:\overline{v}=\sum_{i\notin I}c_{i}\overline{b_{i}}\in\mathbf{N}\otimes\mathbb{R}\ \text{for some}\ I\in\mathcal{A},0\leq c_{i}<1\right\},$
@@
$f]\leftrightarrow v_{f}=\sum_{i=1}^{m}\lceil-(D_{i}\cdot f)\rceil b_{i}.$
@@
where $\iota_{f}=\sum_{i\notin I_{f}}\langle D_{i}\cdot f\rangle$ is the age.
Now assume $P$ arises from a direct sum of line bundles $\oplus_{j=1}^{m}L_{j}$ by removing the zero section, and let $\Lambda_{j}=c_{1}(L_{j}^{*})$. We may take $K_{\mathbb{R}}=(S^{1})^{r}$ and regard the quotient (2.2) as a fiberwise symplectic reduction for the action of $K_{\mathbb{R}}$. This defines $r$ tautological bundles over $\mathcal{P}$. Define $P_{1},\dots,-P_{r}$ to be the $T$-equivariant first Chern classes of the tautological bundles, and $p_{1},\dots,p_{r}$ the restrictions of $P_{1},...,P_{r}$ to the fiber $X_{\omega}$. Recall that $\lambda_{1},\dots,\lambda_{m}$ are the first Chern classes of the $m$ duals of tautological bundles on $BT$ and $R_{T}:=H^{\bullet}_{T}(\text{pt};\mathbb{C})=\mathbb{C}[\lambda_{1},\dots,\lambda_{m}]$. The classes $P_{1},\dots,P_{r}$ generate the cohomology ring $H^{\bullet}_{T}(\mathcal{P})$ over $H^{\bullet}(B)\otimes R_{T}$ just as $p_{1},\dots,p_{r}$ generate $H^{\bullet}_{T}(X_{\omega})$ over $R_{T}$. This gives a basis of $H^{\bullet}_{T}(\mathcal{P})$ over $H^{\bullet}(B)\otimes R_{T}$ consisting of monomials in $\{P_{i}\}$.
@@ 2.4.1. Wall-Crossing for Toric Stack Bundles
By the construction of toric stack bundles (2.2), as the stability parameter $\omega$ crosses a wall, $\mathcal{P}$ changes as the fiber $X_{\omega}$ changes.
As in Section 2.3, consider chambers $C_{\pm}\subset\mathbb{L}\otimes\mathbb{R}$ separated by a wall $W$. Let $\omega_{\pm}\in C_{\pm}$ be stability conditions and let
@@
$\tilde{\mathcal{P}}$$\mathcal{P}_{+}$$\mathcal{P}_{-}$$\mathcal{P}_{0},$$\scriptstyle{f_{+}}$$\scriptstyle{f_{-}}$$\scriptstyle{\varphi}$$\scriptstyle{g_{+}}$$\scriptstyle{g_{-}}$
where each map is induced from a map on toric fibers. Especially, the canonical bundles $K_{\mathcal{P}_{\pm}}$ of $\mathcal{P}_{\pm}$ are restrictions of the same line bundle $L_{0}$ on $\mathcal{P}_{0}$ given by a line bundle on $B$ and the character $\sum_{i=1}^{m}D_{i}$. The condition $\sum_{i=1}^{m}D_{i}\in W$ ensures that $L_{0}$ comes from a $\mathbb{Q}$-Cartier divisor $\mathcal{D}_{0}$ on $\mathcal{P}_{0}$. Then we have
$f^{*}_{+}(K_{\mathcal{P}_{+}})=f^{*}_{+}\circ g^{*}_{+}(\mathcal{D}_{0})=f^{*}_{-}\circ g^{*}_{-}(\mathcal{D}_{0})=f^{*}_{-}(K_{\mathcal{P}_{-}}),$
@@ 3. Crepant Transformation Correspondence
In this Section we prove Theorem 0.1. First, we compute the analytic continuation of the $I$-function using the Mellin-Barnes method, which gives the linear transformation $\mathbb{U}$. Then we compute the Fourier-Mukai transformation on $K$-groups and compare it with $\mathbb{U}$.
@@ 3.1. $I$-Function and $H$-Function
The $I$-function for a toric stack bundle $\mathcal{P}$ is defined in [14] to be
(3.1) $I^{\mathcal{P}}(y,z)=ze^{\sigma/z}\sum_{\mathfrak{D}\in\overline{NE}(B)}J^{B}_{\mathfrak{D}}(z,\tau)Q^{\mathfrak{D}}\sum_{d\in\mathbb{K}}y^{d}\Delta\mathbf{1}_{\left[-d\right]},$
@@
The main result of [14] states that $I^{\mathcal{P}}$ lies on the Lagrangian cone for $\mathcal{P}$. Hence $I^{\mathcal{P}}$ can be used to determine genus 0 Gromov-Witten theory of $\mathcal{P}$. See [4] for the case of toric bundles.
@@
$={z^{-(\sum_{i=1}^{m}(D_{j}\cdot d+\langle D_{j}\cdot-d\rangle)+(\sum_{i=1}^{m}\Lambda_{j})\mathfrak{D})}}=\frac{1}{z^{\sum_{i=1}^{m}D_{j}\cdot d-c_{1}(P)\mathfrak{D}}}\cdot\frac{1}{z^{\iota_{[-d]}}},$
@@
and $\rho=\theta(D_{1}+\cdots+D_{m})$. The notation deg is the degree operator for Chen-Ruan orbifold cohomology and $\text{deg}_{0}$ is the degree operator for ordinary cohomology, i.e. the degree without age shifting. The degrees for variables are given by $\text{deg}(z)=2$ and
@@ 3.2. Analytic Continuation of $H$-function via Localization
For a minimal anticone $\delta\in\mathcal{A}$, there is a corresponding $T$-fixed point in each fiber $X_{\omega}$ over $b\in B$, and together they form a gerbe $B_{\delta}$ over $B$. Let $i_{\delta}:B_{\delta}\rightarrow\mathcal{P}$ be the inclusion map. Similarly, for inertia stack $I\mathcal{P}$, a $T$-fixed gerbe corresponds to a pair $(\delta,f)$ where $\delta\in\mathcal{A}$ and $f\in\mathbb{K}/\mathbb{L}$ satisfy $D_{i}\cdot f\in\mathbb{Z}$ for all $i\in\delta$. Let $i_{(\delta,f)}:B_{(\delta,f)}\rightarrow\mathcal{P}_{f}$ be the inclusion to $I\mathcal{P}$. We have the following restriction of the $H$-function to a fixed gerbe:
@@
For the rest of this Section, any notation with index ++ (respectively -) means it is for $\mathcal{P}_{+}$ (respectively $\mathcal{P}_{-}$). Especially, let $I_{\pm}$ denote their $I$-functions and $H^{\mathfrak{D}}_{\pm}$ denote their $H$-functions respectively. $(\delta_{+},f_{+})$ indexes a $T$-fixed gerbe in $I\mathcal{P}_{+}$ and $(\delta_{-},f_{-})$ indexes a $T$-fixed gerbe in $I\mathcal{P}_{-}$. By $\delta_{+}|\delta_{-}$ we mean $\delta_{+}=\{j_{1},\dots,j_{r-1},j_{+}\},\delta_{-}=\{j_{1},\dots,j_{r-1},j_{-}\}$ with $D_{j_{1}},\dots,D_{j_{r-1}}\in W$, $D_{j_{+}}\cdot e>0,D_{j_{-}}\cdot e<0$. By $(\delta_{+},f_{+})|(\delta_{-},f_{-})$ we mean $\delta_{+}|\delta_{-}$ and there exists $\alpha\in\mathbb{Q}$ such that $f_{-}=f_{+}+\alpha e$.
@@
$i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}=i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{-}.$
@@
$i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}=\sum_{(\delta_{-},f_{-})|(\delta_{+},f_{+})}C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}i^{*}_{(\delta_{-},f_{-})}H^{\mathfrak{D}}_{-},$
@@
$C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}$ $=e^{\frac{\pi iw}{D_{j_{-}}\cdot e}(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}(f_{+}-f_{-}))}$
$\cdot\frac{\sin\pi(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}(f_{+}-f_{-}))}{(-D_{j_{-}}\cdot e)\sin\frac{\pi}{-D_{j_{-}}\cdot e}(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}(f_{+}-f_{-}))}\cdot\prod_{j:D_{j}\cdot e<0,j\neq j_{-}}\frac{\sin\pi(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot f_{+})}{\sin\pi(\frac{U_{j}(\delta_{-})}{2\pi i}+D_{j}\cdot f_{-})},$
@@
Let $\delta_{-}\in\mathcal{A}_{-}$ and $\delta_{+}\in\mathcal{A}_{+}$ be minimal anticones such that $\delta_{-}|\delta_{+}$, and $j_{-}\in\delta_{-}$ the unique element which is not in $\delta_{+}$. Then, for any $j\in[m]$, we have
$U_{j}(\delta_{+})=U_{j}(\delta_{-})+\frac{D_{j}\cdot e}{D_{j_{-}}\cdot e}U_{j_{-}}(\delta_{+}).$
@@
$U_{j}-\lambda_{j}=c_{1}(U_{1}-\lambda_{1})+\cdots+c_{r-1}(U_{r-1}-\lambda_{r-1})+c_{j_{-}}(U_{j_{-}}-\lambda_{j_{-}}).$
@@
$U_{j}(\delta_{-})-\lambda_{j}=\sum_{i=1}^{r-1}c_{i}(-\lambda_{i})+c_{j_{-}}(-\lambda_{j_{-}}),$
$U_{j}(\delta_{+})-\lambda_{j}=\sum_{i=1}^{r-1}c_{i}(-\lambda_{i})+c_{j_{-}}(U_{j_{-}}(\delta_{+})-\lambda_{j_{-}}).$
@@
The above proof shows that this lemma can be generalized to $i^{*}_{\delta_{+}}\theta_{+}(p)=i^{*}_{\delta_{-}}\theta_{-}(p)+\frac{p\cdot e}{D_{j_{-}}\cdot e}U_{j_{-}}(\delta_{+})$ for any $p\in\mathbb{L}\otimes\mathbb{C}$. Especially, we have $\sigma_{+}(\delta)=\sigma_{-}(\delta)$ for $\delta\in\mathcal{A}_{-}\cap\mathcal{A}_{+}$ and $\sigma_{+}(\delta_{+})=\sigma_{-}(\delta_{-})+\frac{(\text{log}(y^{+}))^{e}}{D_{j_{-}}\cdot e}U_{j_{-}}(\delta_{+})$ for $\delta_{+}|\delta_{-}$.
@@
Define $\delta^{\vee}_{+}=\{d\in\mathbb{L}\otimes\mathbb{R}:D_{j}\cdot d+\Lambda_{j}\mathfrak{D}\geq 0,\forall j\in\delta_{+}\}$. Note that the summation in $i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}$ is actually over $d\in\delta^{\vee}_{+}$. Also for $j\in\delta_{+}$, $D_{j}\cdot e=0$ except for one $j_{+}$ where $D_{j_{+}}\cdot e>0$, so for any $d\in\delta^{\vee}_{+}$ we may write $d=d_{+}+ke$ where $d_{+}\in\delta^{\vee}_{+}$ but $d_{+}-e\notin\delta^{\vee}_{+}$, $k\in\mathbb{Z}_{\geq 0}$. Then